• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 16:20:13 on June 26, 2010

    Entry number 328452

    Followups:

    keyword=Day Shift Summary

    June 26, 2010
    SL: Vincent Sulkosky
    TO: Jin Huang
    3rd: N. Toro

    Day Shift Summary:
    ++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++
    Plan: Optics Commissioning
    08:55 Ask for controlled access to insert left HRS sieve-slit.
    09:10 Septum off for access.
    09:22 Move target to Carbon.
    09:23 Turn Septum back on.
    09:28 Ask for 2 uA on carbon.
    09:30 Ask for beam position at BPMB to be (0,0.6).
    09:43 Move to Ta target.
    09:45 Ta target in beam.
    09:55 Move to Optics target.
    10:00 Optics target in, ask for 2 uA.
    10:11 Initial Optics data taken. Turning operations over for detector commissioning.
    10:13 LHRS momentum set to 0.8 GeV/c.
    10:13-11:25 problems getting left dipole NMR to lock.
    11:15 Ask for access to cycle network switch in L-HRS, since all components are not communicating.
    11:45 Dave comes in to begin target cool down.
    13:45 Everyone is out of the hall. But staying in controlled access until target is ready.
    16:00 Target cool down completed.

    Run Summary:
     LHRS Run #   Beam Current   Target   Sieve   Momentum [GeV/c]   Events   T1 Rate   Raster   delta [%]   Comments 
    1126 0.5 uA Optics OUT 1.160 1.0M 800 kHz Off 0. Optics
    1127 0.5 uA Optics OUT 1.160 900k(?) 800 kHz Off 0.Optics
    1128 2 uA Carbon (medium) IN 1.160 500k 31 kHz Off 0.Optics
    1129 1.8 uA Ta IN 1.160 500k 36 kHz Off 0.Optics
    1131 2 uA Optics IN 0.8 N/A N/A Off N/ADAQ crash